Feb 17, 2011 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 10:30 a.m. EST

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • JLab Phone: in CC F326 is 757-269-6337
  • Regina Phone: in LB228 is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or www.readytalk.com (and code without the #)

Items for followup from previous meeting(s)

  1. Monitoring
    • Specifications for BCAL Monitoring R&D (Zisis, Elton, Christina)
    • Zisis will gather Andrei's tests and overall ideas for BCAL monitoring system.
    • IU: checking cross talk of calibration inputs across modules.
  2. Fcal construction
    • Matt to provide optimum schedule for shipping completed blocks to JLab based on technician work plan.
    • Zisis/George to send Elton shipping specifications as example for specifying Pb shipments
    • Mechanical: Tim has promised review of support structure.

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. Calibration
  4. FCAL
    1. Construction Status: mechanical, bases
  5. BCAL
    1. Construction Status: builds, fibre QA
  6. Simulations
    1. Readout segmentation (summing, towers vs quadrants/blocks)
    2. Reconstruction: Study of Segmentation with BCal Reconstruction Code (A. Semenov for UofR group)
  7. Other Updates

Splitoffs discussion:

  • slides (from May 2010 collab. meeting)



Attending: Elton, Yi, tim, John, Beni, David, Tim, Elton (JLab); Zisis, Andrei (UofR); Ryan, Matt (IU); Christina, George (Athens); Will (USM).

  1. Announcements
  2. Review of Action Items
  3. Calibration
    1. Questions about on how to proceed with project
      • Fcal: Matt proposed that a drawing of the FCAL with the plexiglass be put together and sent to Athens (and others) for comments and suggestions for connections to LED sources. Tim can suggest thickness of plexiglass necessary to maintain rigidity.
      • Bcal: Likely path will be using LED sources with connections to fibers which actually get attached to the light guide. Zisis will summarize measurements at UofR.
    2. MOU
      • Christina has sent around an update to their MOU, which includes Stratos Anassontzis, assistant prof, who has agreed to joint the project. He was responsible for the Barrel RICH calibration system of DELPHI with fibers!
  4. FCAL
    1. Mechanical
      • Proceeding well. All good blocks have been prepared. Beginning to use best of "okay" blocks. When these are used up, we will need a procedure for determining formal acceptance of blocks of lesser quality in the detector. Not yet clear how many blocks are actually on hand and how many lower quality blocks will be used.
      • Starting to see if purchase of mu-metal shields can be advanced. These are only needed for assembly and most will likely be dropped shipped to JLab.
    2. Electronics
      • 250 bases have been delivered of the 500 ordered
      • Running into some assembly issues. Initial easy problems have been addressed and fed back to vendor.
      • Remaining 10 bases have more serious issues which are being investigated for patters
      • Paul is planning to visit vendor with feedback before second batch of 250 bases are started. Vendor is very responsive to improving the product
      • The "good" bases are being powered up. At the moment about 100 are connected to the CAN bus and being powered up. Encountering problems with board communication on this large system, but they are being addressed one by one.
    3. Shipments
      • Tim inquired about dates for shipping
      • Likely late March, early April
      • Matt requested information about shipping experience both from Regina and from Jlab.
  5. BCAL
    1. Construction Status: builds, fibre QA
      • Mods 29, 30 stacking complete
      • Mods 31, 32 30% stacked
      • Ross machine shop is delayed by 2 weeks. After startup 8 modules will be machined. Expect shipment of 25-28 end of March.
    2. Fibers
      • Shipment 21 arrived. Testing #19, #18 completed.
      • Retesting previous samples from 5 different shipments to determined degradation of pmt. A small degradation (to be reported later) is observed.
  6. Simulations
    1. Reconstruction: Study of Segmentation with BCal Reconstruction Code (A. Semenov for UofR group)
      • Very non-Gaussian energy deposition shapes observed
      • Selected photons with no conversion before Bcal
      • Generated photons 0-2 GeV, binned in 50 MeV bins.
      • Computed resolution for fine (40 cells) and nominal summed (inner:2x4, outer:2x2)
      • Significant 1/E term observed in the summed cells, with 20-30% worse resolution
      • Significant discussion followed: especially regarding the optimum method for setting thresholds for the summed option.
      • Yi: suggested that for a fair comparison the same number of cells above threshold should be used instead of the same fraction
      • Elton: also the effect of relative reduced fluctuations in a group should be taken into account.
  7. Other Updates

  1. Splitoffs discussion:
    • slides (from May 2010 collab. meeting)
    • Discussion on how to pursue this issue and how it relates to the segmentation. They are clearly issues that affect one another. Discussion will follow at the next calorimeter meeting.